OpenStudy (anonymous):

2(a+3/2)-3(a-2)=2(a-3)

OpenStudy (anonymous):

okay what is the question here??

OpenStudy (anonymous):

take it this is solving for A

OpenStudy (anonymous):

In fact will take each factor and times them for you. 2a=3/2(2)-3a-6+2a-6 is this worked out 2(a+3/2)-3(a-2)=2(a-3) now you take it from here..

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (anonymous):

you still here?